An investor divided $4,000 into two unequal parts. One of them was invested at 5.5% annual simple interest, and the other at 6.7%.
After 1 year, the interest on both investments was $253.
Find how much money was invested at 5.5%.

Answer by richwmiller(17219)

Total amount of money invested: $4000
x+y=4000,
Total yearly interest for the two accounts is: $253
0.055*x+0.067*y=253
x=4000-y
Substitute for x
0.055*(4000-y)+0.067*y=253
Multiply out
220.0-0.055y+0.067*y=253
Combine like terms.
0.012*y=33.0
Isolate y
y=$ 2750.00 at 6.7%
x=4000-y
Calculate x
x=$ 1250.00 at 5.5%
Check
0.055*1250.0+0.067*2750.0=253
68.75+184.25=253
253.0=253
If this statement is TRUE and neither x nor y is negative then all is well